1) A five-year project is expected to generate annual revenues of $159,000, variable costs of $72,500, and fixed costs of $15,000. The annual depreciation is $19,500 and the tax rate is 21 percent. What is the annual operating cash flow?

2) Your local athletic center is planning a $1.2 million expansion to its current facility. This cost will be depreciated on a straight-line basis over a 20-year period. The expanded area is expected to generate $745,000 in additional annual sales. Variable costs are 39* percent of sales, the annual fixed costs are $140,000, and the tax rate is 21 percent. What is the operating cash flow for the first year of this project?
Answer #1

1)

Operating cash flow =(Sales - variables costs - fixed costs - depreciation)(1 - tax) + depreciation

Operating cash flow =(159,000 - 72,500 - 15,000 - 19,500)(1 - 0.21) + 19,500

Operating cash flow = 41,080 + 19,500

Operating cash flow = $60,580

2)

Annual depreciation = 1,200,000 / 20 = 60,000

Variable costs = 0.39 * 745,000 = 290,550

Operating cash flow =(Sales - variables costs - fixed costs - depreciation)(1 - tax) + depreciation

Operating cash flow =(745,000 - 290,550 - 140,000 - 60,000)(1 - 0.21) + 60,000

Operating cash flow = 201,015.5 + 60,000

Operating cash flow = $261,015.50